OpenStudy (anonymous):

write the fraction in LOWEST terms (9x^2/4y^2)

OpenStudy (anonymous):

the fraction is in lowest terms since there are no common factors between 9 and 4; and obviously there are no common factors between x and y \[\frac{9x^2}{4y^2}\] however, since all the factors are pefect squares you can write \[\left( \frac{3x}{2y} \right)^2\]